OpenStudy (anonymous):

ACT math question: A tortoise moves x times as fast as a snail. A hare moves y times as fast as the tortoise. The hare moves how many times as fast as the snail? a)y/x b)x/y c)xy d)y-x E)y+x I tried to plug in numbers but it never seems to work?

OpenStudy (anonymous):

@ganeshie8 @kirbykirby

OpenStudy (kirbykirby):

Let's try a numerical example, and abstract it to numbers Say the snail has a speed of 3 (say a=3) Now, the tortoise moves x times as fast as the snail. Let's say it moves twice as fast. So, x=2. This means that the tortoise moves at a speed 2(3) = x(a) = xa Now, the hare moves y times as fast as the tortoise. So, let's say he moves 4 times as fast (y=4) as the tortoise, meaning 4*(2(3)) = 4*2*3=y*x*a Now... how fast is the hare moving in comparison to the snail? The hare has a speed of 4*2*3 = y*x*a The snail, a speed of 3 = a So how much faster is the hare compared to the snail? can you see the result?

OpenStudy (nikato):

It's actually just a simple substitution problem

OpenStudy (anonymous):

@kirbykirby thanks for replying give me a second to work it out :P

OpenStudy (anonymous):

ok I see it now! I did not think of giving the snail a number.. I substituted the numbers really weirdly. thanks!

OpenStudy (kirbykirby):

:)

OpenStudy (nikato):

Here's another way. It'll be less time consuming if you understand it, or you can continue solving by plugging numbers. Tortoise moves x times as fast as snail T= Sx Hare moves y times as fast as tortoise. H= Ty Substitute the first equation into second equation. H= (Sx)y So H= Sxy So hare is xy times as fast as snail